A solid cylinder has a radius of 5cm and a height of 12cm. Calculate the total surface area.
Is this how you work it out:
3.142 x 5² = 78.55
78.55 x 12 = 942.6 ?

Answer by edjones(7569)

lateral surface area=pi*2r*h
=pi*10*12
=120*pi
.
Top and bottom s.a.=2*pi*r^2
=pi*25*2
=50*pi
.
120pi+50pi
=170pi or 534 cm
.
Ed